# Chao-Han Huck Yang1, Ramani Duraiswami2, Dinesh Manocha2, Rafael Valle1, Bryan Catanzaro1

### v1

@@ 1 Introduction
Audio-Language Models (ALMs) extend the capabilities of LMs to the auditory domain. Early works such as CLAP elizalde2022clap align audio and text in a shared embedding space, enabling them with tasks like text-to-audio retrieval (oncescu2021audio, ). More recently, the emergence of Large ALMs (LALMs)—decoder-only language models augmented with audio understanding chu2023qwenaudio ; chu2024qwenaudio2 ; team2023gemini —has unlocked powerful capabilities, including open-ended audio question-answering (AQA) that demands both reasoning and world knowledge (sakshi2024mmau, ). These capabilities have further enabled tasks like audio analysis drossos2020clotho ; kim2019audiocaps , conversational assistants daniel2018toward , etc.
@@
However, existing models still fall short in key areas critical to AGI, such as expert-level reasoning morris2024position ; sakshi2024mmau , multi-turn and multi-audio dialogue goel2024audio , and long audio understanding kong2025audioflamingo2 . We identify two core limitations: (i) most LALMs are trained primarily on short audio for recognition tasks rather than ones that require deliberate reasoning; and (ii) in turn, they lack exposure to the skill sets required for complex tasks. Additionally, most LALMs that support all three modalities of speech, sound, and music are closed-source: while some have publicly released models weights chu2023qwenaudio ; chu2024qwenaudio2 ; abouelenin2025phi , they offer limited to no information about their data, code, or recipes (more details in Table 1).
Main Contributions. To address these issues, we introduce Audio Flamingo 3 (AF3), a fully open-source LALM with state-of-the-art performance in audio understanding and reasoning across 20+ benchmarks. In addition, AF3 brings several novel capabilities, including multi-turn, multi-audio chat, on-demand thinking, voice-to-voice interaction, and long-context audio reasoning (up to 10 minutes). We propose three core innovations to enable these capabilities: (i) Data: We focus on curating high-quality data at scale and propose (a) AudioSkills-XL: a large-scale dataset of 8M diverse AQA pairs, (b) LongAudio-XL: large-scale dataset of 1.25M diverse audio QA pairs for long audio reasoning; (c) AF-Chat: a multi-turn multi-audio chat dataset curated using a novel algorithm with 75k instances and (d) AF-Think: a dataset with 250k+ AQA pairs with short length prefixes to encourage CoT-type reasoning before arriving at the answer (ii) AF-Whisper: We train AF-Whisper, a unified audio encoder pretrained using a novel strategy on large-scale audio-caption pairs, capable of learning general-purpose representations across speech, sounds, and music; and (iii) Learning Curriculum: We train AF3 with a five-stage curriculum-based training strategy that progressively increases context length and task complexity. In summary, our main contributions are:
@@ 3.1 Audio Flamingo 3 Architecture
In this section, we discuss our proposed architecture for Audio Flamingo 3 as shown in Figure 2. AF3 consists of i) AF-Whisper: an audio encoder with sliding window feature extraction, ii) audio projector, iii) an LLM, and iv) a streaming TTS. We provide details of each component below.
@@
As illustrated in Figure 2, we start with the pre-trained Whisper large-v3 encoder radford2022whisper , attach it to a standard Transformer decoder, and train using the audio captioning task with the next-token-prediction objective. To achieve this, we generate a natural language caption for each audio, describing its speech, sound, and music content. First, we pool several datasets and then prompt GPT-4.1 to generate the audio caption. For prompting, we use available metadata for each sample, which includes transcripts, ambient sound descriptions, and music attributes. For samples lacking any of the 3 metadata, we synthesize it using AF2 (kong2025audioflamingo2, ) or Whisper-Large-v3 ASR (radford2022whisper, ). All datasets used for training are detailed in Section A.2. We choose Whisper as the backbone due to its existing speech understanding capabilities and its dense, high-resolution audio features, which are more informative than those from models like CLAP (elizalde2022clap, ). We connect it with a Transformer decoder using cross-attention (similar to RECAPliu2023recap and AF2 (kong2025audioflamingo2, )) with 24 layers, 8 attention heads, and 1024 hidden size.
Feature Extraction. Given an audio input $A$, we first resample it to 16kHz mono. The raw waveform is then transformed into a 128-channel mel-spectrogram using a window size of 25ms and a hop size of 10ms. This mel-spectrogram is processed by AF-Whisper, producing hidden representations, denoted as $h_{a}=f_{a}(A)$, where $h_{a}\in\mathbb{R}^{N\times d}$. As shown in Figure 2, each audio is processed in 30-second chunks of non-overlapping sliding windows, and $N$ or the temporal resolution depends on the length of the audio and the maximum number of sliding windows (which varies according to the stage of training). AF-Whisper produces audio features at a frame rate of 50Hz, and we further apply a pooling layer with a stride of two similar to chu2024qwenaudio2 . $d$ denotes the hidden dimension, which is 1280.
@@ 4.2 LongAudio-XL: Expanding LongAudio with Long Speech QA
To our knowledge, Long Speech QA (i.e., audio $\geq$ 30 seconds) has not been explored in prior work, despite its relevance to real-world applications such as long-form conversation understanding, meeting summarization, and narrative comprehension. To bridge this gap, we extend the existing LongAudio dataset (kong2025audioflamingo2, ) (focused on sounds and music) by incorporating over 1M reasoning-focused QA examples from long-form speech (30s-10min). We curate audios from diverse sources including: Single-speaker speech: LibriSpeech (audiobooks) (panayotov2015librispeech, ), EuroParl (koehn2005europarl, ), VoxPopuli (parliamentary debates) (wang2021voxpopuli, ) and Multi-speaker conversations: Spotify Podcasts (clifton-etal-2020-100000, ), Switchboard (godfrey1992switchboard, ), Fisher (dyadic calls) (cieri2004fisher, ), MELD (poria2018meld, ), DailyTalk (lee2023dailytalk, ), MMDialog (natural dialogues) (feng2022mmdialog, ). We merge consecutive short segments in chronological order to construct longer, coherent audios. We construct QAs across a wide range of skills, as illustrated in Figure 3:
@@ 4.3 AF-Think: Towards flexible, on-demand reasoning
In AF3, we adopt a lightweight thinking mechanism with two key modifications: (i) We create AF-Think, a dataset of 250k MCQ-based QAs with short, controlled thought preceding the answer. This additional thinking serve as a prefix to the answer and are limited to an average of approximately 40 words, providing concise yet effective context for audio QA (example in Figure 3). (ii) Instead of explicitly post-training for CoT, we add a special suffix to QA prompts (highlighted in Figure 3). We include AF-Think in the Stage 3.5 training mixture, upweighted relative to standard QA data. This allows AF3 to think only when prompted, offering flexible, on-demand additional reasoning.
@@ 4.4 AF-Chat: Multi-turn Multi-audio Chat Data
Next, we prompt GPT-4.1 using carefully designed expert exemplars (Fig. 35 and  34) to generate natural, multi-turn chat sessions under the following constraints: (i) the model may choose any subset of the similar/dissimilar audios (up to 10 turns), prioritizing conversation quality; (ii) not all turns require a new audio—follow-up and clarification questions are encouraged; and (iii) later turns may refer back to earlier audios or responses to simulate real conversational grounding. The design of AF-Chat is informed by extensive internal human studies to reflect how users naturally interact with audio-language models. As a result, it provides rich, diverse supervision for aligning LALMs to handle complex, contextual, and naturalistic audio conversations. Finally, we select 200 high-quality samples for the test set, known as AF-Chat-test, and ensure that the audios in these instances have audio clips that were not seen during training.
@@ 6 Experiments
Audio Flamingo 3 76.67 | 73.30 | 64.87 | 72.28
+Think 76.85 | 73.92 | 66.07 | 73.14
@@
Experimental Setup. We train AF3 on 128 NVIDIA A100 GPUs, each with 80GB of memory. Details about batch size, learning rates, and optimizers for each stage of training are in Appendix H.
@@
Evaluation Datasets. We evaluate AF3 on a variety of tasks and benchmarks, including audio classification (CochlScene (jeong2022cochlscene, ), NSynth (Source and Instrument) (engel2017neural, ), NonSpeech7k (rashid2023nonspeech7k, ), IEMOCAP (busso2008iemocap, )), audio QA (ClothoAQA (lipping2022clotho, ), MusicAVQA (li2022learning, ), Music Instruct (deng2023musilingo, ), LibriSQA (zhao2023librisqa, )), reasoning-focused audio QA (MMAU (sakshi2024mmau, ) (v05.15.25), MuchoMusic (perceptual version) (zang2025you, ; weck2024muchomusic, ), MMAR (ma2025mmarchallengingbenchmarkdeep, ), MMSU (wang2025mmsu, ), CompA-R-test (ghoshcompa, ), Audio Entailment (deshmukh2025audio, )), multimodal hallucination detection (CMM (leng2024curse, )), audio captioning (Clotho-v2 (drossos2020clotho, ), AudioCaps (kim2019audiocaps, )), ASR (Librispeech (clean and other) (panayotov2015librispeech, ), SPGISpeech (o2021spgispeech, ), TEDLIUM (rousseau2012ted, ; hernandez2018ted, ), GigaSpeech (Large) (chen2021gigaspeech, ), Common Voice 15 (commonvoice:2020, ) and Voxpopuli (wang2021voxpopuli, )) and long audio captioning and QA (LongAudioBench – which we augment with 2.5k human-annotated long-speech QA instances). For evaluating chat capabilities, we conduct a human study of model outputs on AF-Chat-test (more details in Appendix E) and compare only with Qwen2-Audio. Each annotator is asked to rate the response of the model for every turn on a scale of 1-5 for factuality, usefulness, and depth. We report results averaged across all instances across all turns. Furthermore, we evaluate the voice-text capabilities of our AF3-Chat model on two datasets, OpenAudioBench (li2025baichuan, ) and VoiceBench (chen2024voicebench, ). These benchmarks consist of voice queries (synthetically generated speech from text queries) and assess aspects such as instruction following, question answering, trivia knowledge, and reasoning. Finally, we evaluate our speech generation module using zero-shot TTS evaluation on the English subset of the SEED benchmark (anastassiou2024seed, ). To calculate accuracy, we use either exact string matching with the ground truth or CLAP-based retrieval following (deshmukh2023pengi, ), implemented with open-source AF-CLAP (kong2025audioflamingo2, ). For MCQ, AF3 typically outputs only the selected option. In cases where the model provides more verbose or open-ended responses (e.g., with thinking mode), we apply multiple regex patterns to extract the chosen option.
@@ 6.1 Audio Understanding and Reasoning Evaluation
AF3 is the strongest and fully open-source LALM. Table 2 shows AF3 outperforming previous SOTA open-weight and closed-source models across a wide range of audio understanding and reasoning benchmarks. AF3 sets new highs on MMAU (72.28) (note for Qwen2.5-Omni on MMAU we report the “parsed score” for fair evaluation), ClothoAQA (91.1), Clotho Entailment (92.9), and CMM Hallucination (86.7). On tasks like NSynth and MusicInstruct, it shows significant gains, highlighting strong sound and music understanding. For LongAudioBench (sound and speech), AF3 outperforms Gemini 2.5 Pro by a wide margin, demonstrating its strength in long-context reasoning. We also evaluate AF3 with thinking prompts (+Think) on reasoning-heavy benchmarks like MMAU and MuchoMusic, observing a performance boost. Additionally, AF3 achieves state-of-the-art ASR results on LibriSpeech, SPGISpeech, and VoxPopuli—even compared to dedicated ASR models—despite not being trained on large-scale ASR datasets like many open-weight models.
@@ 6.2 Chat and TTS Evaluation
Voice-Text and Speech Generation Evaluation. Table 3 evaluates AF3-Chat on two key tasks: voice-to-text and text-to-speech generation. In the voice-to-text setting (spoken QA), AF3-Chat achieves strong gains across all of OpenAudioBench, surpassing Qwen2.5-Omni. On VoiceBench, which tests spoken QA robustness across AdvBench, CommonEval, and OpenBookQA, AF3-Chat performs comparably to Qwen2.5-Omni and Qwen2-Audio Chat. For TTS (evaluated on SEED test-en), AF3-Chat shows improved performance with a lower WER of 2.02 (vs. 2.72 for Qwen2.5-Omni) and a speaker similarity score of 0.61, closely matching Qwen2.5’s 0.63.
@@ 6.3 Ablation Studies
In this section, we ablate our key components (using just 10% of the training data) to support the paper’s main claims. More ablations are provided in Appendix J.
Evaluating AF-Whisper as a Unified Encoder. Table 4 compares AF3 trained with our unified AF-Whisper encoder against a dual-encoder setup using CLAP for sounds/music and Whisper-v3 for speech (elizalde2022clap, ; radford2022whisper, ). AF-Whisper outperforms the dual-encoder model under the same data budget, demonstrating its effectiveness as a single encoder for sound, music, and speech.
AudioSkills-XL: A Key Dataset for Performance Gains.: To measure the impact of AudioSkills-XL, we ablate it from Stage 3 of training and compare results to the full setup. As shown in Table 4, removing AudioSkills-XL causes a significant performance drop—particularly on MMAU—underscoring its role in improving generalization and robustness. These findings highlight the value of large-scale, skill-targeted audio QA data for fine-tuning multi-modal models.

### v2

@@ 1 Introduction
Audio-Language Models (ALMs) extend the capabilities of LMs to the auditory domain. Early works such as CLAP elizalde2022clap align audio and text in a shared embedding space, enabling them with tasks like retrieval (oncescu2021audio, ). More recently, the emergence of Large ALMs (LALMs)—decoder-only language models augmented with audio understanding chu2023qwenaudio ; chu2024qwenaudio2 ; team2023gemini —has unlocked powerful capabilities, including open-ended audio question-answering (AQA) that demands both reasoning and world knowledge (sakshi2024mmau, ). These capabilities have further enabled tasks like audio analysis drossos2020clotho ; kim2019audiocaps , conversational assistants daniel2018toward , etc.
@@
However, existing models still fall short in key areas critical to AGI, such as expert-level reasoning morris2024position ; sakshi2024mmau , multi-turn and multi-audio dialogue goel2024audio , and long audio understanding kong2025audioflamingo2 . We identify two core limitations: (i) most LALMs are trained primarily on short audio for recognition tasks rather than ones that require deliberate reasoning; and (ii) in turn, they lack exposure to the skill sets required for complex tasks. Additionally, most LALMs that support all three modalities of speech, sound, and music are closed-source: while some have publicly released models weights chu2023qwenaudio ; chu2024qwenaudio2 ; abouelenin2025phi , they offer limited to no information about their data, code, or recipes (more details in Table˜1).
Main Contributions. To address these issues, we introduce Audio Flamingo 3 (AF3), a fully open-source [^1] By fully open, we mean that the model’s weights, training data, and code will be publicly released, with full transparency about the training methodology. Due to the licensing and scope of the training data used in the work, all releases will be under a research-only license. LALM with state-of-the-art performance in audio understanding and reasoning across 20+ benchmarks. In addition, AF3 brings several novel capabilities, including multi-turn, multi-audio chat, on-demand thinking, voice-to-voice interaction, and long-context audio reasoning (up to 10 minutes). We propose three core innovations to enable these capabilities: (i) Data: We focus on curating high-quality data at scale and propose (a) AudioSkills-XL: a large-scale dataset of 8M diverse AQA pairs, (b) LongAudio-XL: large-scale dataset of 1.25M diverse audio QA pairs for long audio reasoning; (c) AF-Chat: a multi-turn multi-audio chat dataset curated using a novel algorithm with 75k instances and (d) AF-Think: a dataset with 250k+ AQA pairs with short length prefixes to encourage CoT-type reasoning before arriving at the answer (ii) AF-Whisper: We train AF-Whisper, a unified audio encoder pretrained using a novel strategy on large-scale audio-caption pairs, capable of learning general-purpose representations across speech, sounds, and music; and (iii) Learning Curriculum: We train AF3 with a five-stage curriculum-based training strategy that progressively increases context length and task complexity. In summary, our main contributions are:
@@ 3.1 Audio Flamingo 3 Architecture
In this section, we discuss our proposed architecture for Audio Flamingo 3 as shown in Figure˜2. AF3 consists of i) AF-Whisper: an audio encoder with sliding window feature extraction, ii) audio projector, iii) an LLM, and iv) a streaming TTS. We provide details of each component below.
@@
As illustrated in Figure˜2, we start with the pre-trained Whisper large-v3 encoder radford2022whisper , attach it to a standard Transformer decoder, and train using the audio captioning task with the next-token-prediction objective. To achieve this, we generate a natural language caption for each audio, describing its speech, sound, and music content. First, we pool several datasets and then prompt GPT-4.1 to generate the audio caption. For prompting, we use available metadata for each sample, which includes transcripts, ambient sound descriptions, and music attributes. For samples lacking any of the 3 metadata, we synthesize it using AF2 (kong2025audioflamingo2, ) or Whisper-Large-v3 ASR (radford2022whisper, ). All datasets used for training are detailed in Section A.2. We choose Whisper as the backbone due to its existing speech understanding capabilities and its dense, high-resolution audio features, which are more informative than those from models like CLAP (elizalde2022clap, ). We connect it with a Transformer decoder using cross-attention (similar to RECAPliu2023recap and AF2 (kong2025audioflamingo2, )) with 24 layers, 8 attention heads, and 1024 hidden size.
Feature Extraction. Given an audio input $A$, we first resample it to 16kHz mono. The raw waveform is then transformed into a 128-channel mel-spectrogram using a window size of 25ms and a hop size of 10ms. This mel-spectrogram is processed by AF-Whisper, producing hidden representations, denoted as $h_{a}=f_{a}(A)$, where $h_{a}\in\mathbb{R}^{N\times d}$. As shown in Figure˜2, each audio is processed in 30-second chunks of non-overlapping sliding windows, and $N$ or the temporal resolution depends on the length of the audio and the maximum number of sliding windows (which varies according to the stage of training). AF-Whisper produces audio features at a frame rate of 50Hz, and we further apply a pooling layer with a stride of two similar to chu2024qwenaudio2 . $d$ denotes the hidden dimension, which is 1280.
@@ 4.2 LongAudio-XL: Expanding LongAudio with Long Speech QA
To our knowledge, Long Speech QA (i.e., audio $\geq$ 30 seconds) has not been explored in prior work, despite its relevance to real-world applications such as long-form conversation understanding, meeting summarization, and narrative comprehension. To bridge this gap, we extend the existing LongAudio dataset (kong2025audioflamingo2, ) (focused on sounds and music) by incorporating over 1M reasoning-focused QA examples from long-form speech (30s-10min). We curate audios from diverse sources including: Single-speaker speech: LibriSpeech (audiobooks) (panayotov2015librispeech, ), EuroParl (koehn2005europarl, ), VoxPopuli (parliamentary debates) (wang2021voxpopuli, ) and Multi-speaker conversations: Spotify Podcasts (clifton-etal-2020-100000, ), Switchboard (godfrey1992switchboard, ), Fisher (dyadic calls) (cieri2004fisher, ), MELD (poria2018meld, ), DailyTalk (lee2023dailytalk, ), MMDialog (natural dialogues) (feng2022mmdialog, ). We merge consecutive short segments in chronological order to construct longer, coherent audios. We construct QAs across a wide range of skills, as illustrated in Figure˜3:
@@ 4.3 AF-Think: Towards flexible, on-demand reasoning
In AF3, we adopt a lightweight thinking mechanism with two key modifications: (i) We create AF-Think, a dataset of 250k MCQ-based QAs with short, controlled thought preceding the answer. This additional thinking serve as a prefix to the answer and are limited to an average of approximately 40 words, providing concise yet effective context for audio QA (example in Figure˜3). (ii) Instead of explicitly post-training for CoT, we add a special suffix to QA prompts (highlighted in Figure˜3). We include AF-Think in the Stage 3.5 training mixture, upweighted relative to standard QA data. This allows AF3 to think only when prompted, offering flexible, on-demand additional reasoning.
@@ 4.4 AF-Chat: Multi-turn Multi-audio Chat Data
Next, we prompt GPT-4.1 using carefully designed expert exemplars (Fig. 36 and  35) to generate natural, multi-turn chat sessions under the following constraints: (i) the model may choose any subset of the similar/dissimilar audios (up to 10 turns), prioritizing conversation quality; (ii) not all turns require a new audio—follow-up and clarification questions are encouraged; and (iii) later turns may refer back to earlier audios or responses to simulate real conversational grounding. The design of AF-Chat is informed by extensive internal human studies to reflect how users naturally interact with audio-language models. As a result, it provides rich, diverse supervision for aligning LALMs to handle complex, contextual, and naturalistic audio conversations. Finally, we select 200 high-quality samples for the test set, known as AF-Chat-test, and ensure that the audios in these instances have audio clips that were not seen during training.
@@ 6 Experiments
Audio Flamingo 3 75.83 | 74.47 | 66.97 | 72.42
+Think 75.27 | 74.60 | 69.60 | 73.16
MMAU-v05.15.25 (test-mini) Sound | Music | Speech | Avg Qwen2.5-O ACC ↑ 78.10 | 65.90 | 70.60 | 71.50
Audio Flamingo 3 79.58 | 73.95 | 66.37 | 73.30
+Think 79.88 | 76.55 | 66.37 | 74.26
@@
Experimental Setup. We train AF3 on 128 NVIDIA A100 GPUs, each with 80GB of memory. Details about batch size, learning rates, and optimizers for each stage of training are in Appendix˜H.
@@
Evaluation Datasets. We evaluate AF3 on a variety of tasks and benchmarks, including audio classification (CochlScene (jeong2022cochlscene, ), NSynth (Source and Instrument) (engel2017neural, ), NonSpeech7k (rashid2023nonspeech7k, ), IEMOCAP (busso2008iemocap, )), audio QA (ClothoAQA (lipping2022clotho, ), MusicAVQA (li2022learning, ), Music Instruct (deng2023musilingo, ), LibriSQA (zhao2023librisqa, )), reasoning-focused audio QA (MMAU (sakshi2024mmau, ) (v05.15.25), MuchoMusic (perceptual version) (zang2025you, ; weck2024muchomusic, ), MMAR (ma2025mmarchallengingbenchmarkdeep, ), MMSU (wang2025mmsu, ), CompA-R-test (ghoshcompa, ), Audio Entailment (deshmukh2025audio, )), multimodal hallucination detection (CMM (leng2024curse, )), audio captioning (Clotho-v2 (drossos2020clotho, ), AudioCaps (kim2019audiocaps, )), ASR (Librispeech (clean and other) (panayotov2015librispeech, ), SPGISpeech (o2021spgispeech, ), TEDLIUM (rousseau2012ted, ; hernandez2018ted, ), GigaSpeech (Large) (chen2021gigaspeech, ), Common Voice 15 (commonvoice:2020, ) and Voxpopuli (wang2021voxpopuli, )) and long audio captioning and QA (LongAudioBench – which we augment with 2.5k human-annotated long-speech QA instances). For evaluating chat capabilities, we conduct a human study of model outputs on AF-Chat-test (more details in Appendix˜E) and compare only with Qwen2-Audio. Each annotator is asked to rate the response of the model for every turn on a scale of 1-5 for factuality, usefulness, and depth. We report results averaged across all instances across all turns. Furthermore, we evaluate the voice-text capabilities of our AF3-Chat model on two datasets, OpenAudioBench (li2025baichuan, ) and VoiceBench (chen2024voicebench, ). These benchmarks consist of voice queries (synthetically generated speech from text queries) and assess aspects such as instruction following, question answering, trivia knowledge, and reasoning. Finally, we evaluate our speech generation module using zero-shot TTS evaluation on the English subset of the SEED benchmark (anastassiou2024seed, ). All baseline results reported in this work are based on our own evaluations; we did not rely on results from prior literature (in some cases, we were unable to reproduce the numbers as originally reported). To calculate accuracy, we use either exact string matching with the ground truth or CLAP-based retrieval following (deshmukh2023pengi, ), implemented with open-source AF-CLAP (kong2025audioflamingo2, ). For MCQ, AF3 typically outputs only the selected option. In cases where the model provides more verbose or open-ended responses (e.g., with thinking mode), we apply multiple regex patterns to extract the chosen option.
@@ 6.1 Audio Understanding and Reasoning Evaluation
AF3 is the strongest and fully open-source LALM. Table˜2 shows AF3 outperforming previous SOTA open-weight and closed-source models across a wide range of audio understanding and reasoning benchmarks. AF3 sets new highs on MMAU (72.42) (note for Qwen2.5-Omni on MMAU we report the “parsed score” for fair evaluation), ClothoAQA (91.1), Clotho Entailment (92.9), and CMM Hallucination (86.7). On tasks like NSynth and MusicInstruct, it shows significant gains, highlighting strong sound and music understanding. For LongAudioBench (sound and speech), AF3 outperforms Gemini 2.5 Pro by a wide margin, demonstrating its strength in long-context reasoning. We also evaluate AF3 with thinking prompts (+Think) on reasoning-heavy benchmarks like MMAU and MuchoMusic, observing a performance boost. Although the thinking mode is activated after Stage 3.5 only when using our specific thinking prompt, the checkpoint remains usable without it. We report average scores of 73.16 and 74.26 on MMAU-test and MMAU-test-mini, respectively. Additionally, AF3 achieves state-of-the-art ASR results on LibriSpeech, SPGISpeech, and VoxPopuli—even compared to dedicated ASR models—despite not being trained on large-scale ASR datasets like many open-weight models. We illustrate a demo of AF3’s capabilities in Fig. 14.
@@ 6.2 Chat and TTS Evaluation
Voice-Text and Speech Generation Evaluation. Table˜3 evaluates AF3-Chat on two key tasks: voice-to-text and text-to-speech generation. In the voice-to-text setting (spoken QA), AF3-Chat achieves strong gains across all of OpenAudioBench, surpassing Qwen2.5-Omni. On VoiceBench, which tests spoken QA robustness across AdvBench, CommonEval, and OpenBookQA, AF3-Chat performs comparably to Qwen2.5-Omni and Qwen2-Audio Chat. For TTS (evaluated on SEED test-en), AF3-Chat shows improved performance with a lower WER of 2.02 (vs. 2.72 for Qwen2.5-Omni) and a speaker similarity score of 0.61, closely matching Qwen2.5’s 0.63.
@@ 6.3 Ablation Studies
In this section, we ablate our key components (using just 10% of the training data) to support the paper’s main claims.
Evaluating AF-Whisper as a Unified Encoder. Table˜4 compares AF3 trained with our unified AF-Whisper encoder against a dual-encoder setup using CLAP for sounds/music and Whisper-v3 for speech (elizalde2022clap, ; radford2022whisper, ). AF-Whisper outperforms the dual-encoder model under the same data budget, demonstrating its effectiveness as a single encoder for sound, music, and speech.
AudioSkills-XL: A Key Dataset for Performance Gains.: To measure the impact of AudioSkills-XL, we ablate it from Stage 3 of training and compare results to the full setup. As shown in Table˜4, removing AudioSkills-XL causes a significant performance drop—particularly on MMAU—underscoring its role in improving generalization and robustness. These findings highlight the value of large-scale, skill-targeted audio QA data for fine-tuning multi-modal models.